OpenStudy (anonymous):

Calculate the discriminant and use it to determine how many real-number roots the equation has. 3x^2 – 6x + 1 = 0 three real-number roots two real-number roots one real-number root no real-number roots

OpenStudy (anonymous):

find \(b^2-4ac\) with \(a=3,b=-6,c=1\)

OpenStudy (anonymous):

one real-number root?

OpenStudy (anonymous):

lets see \[(-6)^2-4\times 3\times 1=36-12=24\] a positive number so two real roots